mathslover (mathslover):

Starting Differentiation... Problem (1) Evaluate \(\cfrac{d}{dx} \left[\cfrac{x^2-\sqrt{x}}{x^2 + 2x} \right]\)

OpenStudy (anonymous):

Quotient rule, know of it?

mathslover (mathslover):

Hmm yeah, I think so...

OpenStudy (goformit100):

Sure the quotient rule will work

mathslover (mathslover):

\(\cfrac{f(x) g'(x) - g(x) f(x) }{(g(x))^2}\)

OpenStudy (anonymous):

Missed a prime, the g'(x) like that on -f(x)g'(x)

mathslover (mathslover):

Hmm... Okay!!

OpenStudy (anonymous):

\[\left( \frac{ f(x) }{ g(x) } \right)' =~ \frac{ g(x)f(x)'-f(x)g'(x) }{ [g(x)]^2 }\]

mathslover (mathslover):

So, it becomes : \(\cfrac{(x^2-\sqrt{x})(2x + 2 ) - (2x - \cfrac{1}{2\sqrt{x}}) }{(x^2 2x)^2}\)

mathslover (mathslover):

Sorry

mathslover (mathslover):

\(\cfrac{(x^2-\sqrt{x})(2x + 2 ) - (2x - \cfrac{1}{2\sqrt{x}})(x^2 + 2x) }{(x^2 + 2x)^2}\)

OpenStudy (anonymous):

It's alright, just take it one step at a time ^.^

mathslover (mathslover):

I think the numerator is messed up... it should be : b -a instead of a - b right?

OpenStudy (anonymous):

\[\huge \frac{ (x^2+2x)(2x-\frac{ 1 }{ 2\sqrt{x} })-(x^2-\sqrt{x})(2x+2) }{ (x^2+2x )^2}\]

ganeshie8 (ganeshie8):

use product rule if quotient rule gets messy...

mathslover (mathslover):

Hmm yeah Batman, I meant that with a-b and b - a :P @ganeshie8 -> can you please elaborate ? just a little bit..

ganeshie8 (ganeshie8):

\[\large \cfrac{d}{dx} \left[\cfrac{x^2-\sqrt{x}}{x^2 + 2x} \right] = \cfrac{d}{dx} \left[(x^2-\sqrt{x})(x^2 + 2x)^{-1}\right] \]

mathslover (mathslover):

Okay... got it!

OpenStudy (anonymous):

Yeah product rule works to lol, should've just done that, \[\left( f(x)g(x) \right)=f'(x)g(x)+f(x)g'(x)\] \[\frac{ d }{ dx }\left[ (x^2-\sqrt{x})(x^2+2x)^{-1} \right]\]

ganeshie8 (ganeshie8):

product rule also gives u messy expressions - but atleast to work it u memorize a simple formula...

mathslover (mathslover):

Yeah.... \(\left[(x^2-\sqrt{x})(-(x^2+2x)^{-2}) *(2x+2)\right] + (2x - \cfrac{1}{\sqrt{x}})(x^2+2x)^{-1}\)

OpenStudy (anonymous):

\[\huge \frac{ 4x ^{\frac{ 3 }{ 2 }}+3x+2 }{ 2x ^{\frac{ 3 }{ 2 }}(x+2)^2 }\] This is what the final answer is as I got, just check if you get that as well, I'm off to bed, take care and gl @mathslover

mathslover (mathslover):

\(\cfrac{(x^2-\sqrt{x})(2x+2)}{-(x^2+2x)^2} + \cfrac{\left(2x-\cfrac{1}{\sqrt{x}}\right)}{(x^2+2x)}\) Am I wrong somewhere?

Miracrown (miracrown):

yes everywhere ! :P

mathslover (mathslover):

Okay, @iambatman - I will check and thanks for your time. I appreciate it a lot. Have a sound sleep.

mathslover (mathslover):

Miracrown, can you please point it out?

ganeshie8 (ganeshie8):

\[\left[(x^2-\sqrt{x})(-(x^2+2x)^{-2}) *(2x+2)\right] + (2x - \cfrac{1}{\color{red}{2}\sqrt{x}})(x^2+2x)^{-1} \]

Miracrown (miracrown):

lol just kiddin... @ganeshie8 can check it for you, I can't be stuffed. Super tired !

ganeshie8 (ganeshie8):

^^ that 2 should be there as derivative of sqrt(x) is 1/2sqrt(x)

mathslover (mathslover):

Oh, fine. @Miracrown Thanks. Oh... yeah, how I missed that \(\color{blue}{\bf 2}\) there... :( Thanks for poiting that out Ganeshie8

ganeshie8 (ganeshie8):

we're done wid differentiating - calculus part is over :) you can simplify the answer using algebra tricks if u want... or leave the answer as it is...

Miracrown (miracrown):

@mathslover show us ur algebra tricks! :P

mathslover (mathslover):

Well , when algebra comes into consideration, I am generally excited about that... I'll try to simply this and post it here ... may be in 2-3 minutes.

Miracrown (miracrown):

Take your time. =)

ganeshie8 (ganeshie8):

its wolfram's job to verify the answer and make u feel proud :) once u have the final answer, click below : http://www.wolframalpha.com/input/?i=%28%28x%5E2-sqrt%28x%29%29%2F%28x%5E2%2B2x%29%29%27

mathslover (mathslover):

I just got a sweet expression...

ganeshie8 (ganeshie8):

Congrats ! :)

mathslover (mathslover):

Unfortunately, I can not post the whole process here @Miracrown , so , sorry, I can't exhibit my skills (algebraic) right now :( ... @ganeshie8 @iambatman @Miracrown -> Thanks for all your help and participation in this post.

Miracrown (miracrown):

That's ok-maybe next time. :-]

mathslover (mathslover):

Yeah, that *next time* is coming just after 5-6 minutes... !

Miracrown (miracrown):

Why now?- I'm super tired, so idk if I'll be much help at this state.

mathslover (mathslover):

I have some more questions, though, I already have tried them 4-5 times, but I will try one more time, and if I don't get them this time also.. I will post them...! When do you think you will be available for help?

Miracrown (miracrown):

Actually bring it on! :)

Miracrown (miracrown):

I'll try to help you with one for now, but the rest maybe next time or someone else will surely give you a hand.

mathslover (mathslover):

Sure.